OpenStudy (anonymous):

what are the solutions of the equation: 0= x^3 + 3x - 10

OpenStudy (anonymous):

x = 5, 2 x= -5, -2 x= -5, 2 x= 5, -2

OpenStudy (anonymous):

i am going to make a guess that it is \[0=x^2-3x-10\] am i right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no wait its +3x

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

do you know the answer @Blank 

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

Could you take a pic cause I can't seem to get the correct answer. I always keep getting: x = 1.6989. I'm still working on it.

OpenStudy (anonymous):

oh im sorry its actually 0= x^2 + 3x - 10

OpenStudy (anonymous):

i accidentally put x^3, sorry!

OpenStudy (blank ):

There you go. :) x=-5,2

OpenStudy (anonymous):

thanks once again (:

OpenStudy (blank ):

You gotta be careful when posting questions or you might get a wrong answer. ǝɯoɔlǝʍ ǝɹ,noʎ :)